Given two independent random samples with the following results: n1=510x1=231   n2=728x2=139 Use this data to find the 98% confidence interval for the true difference between the population proportions. Step 1 of 4 : Find the values of the two sample proportions, pˆ1 and pˆ2. Round your answers to three decimal places
Answer #1

Solution :

From given that,

Step 1:

= x1 / n1 = 231 / 510 = 0.453

= x2 / n2 = 139 / 728 = 0.191